OpenStudy (anonymous):

Find the perimeter and area of a right triangle if one leg measures 10 ft. and the other leg measures 24 ft. must show all work step by step . Include correct units with your solution.

OpenStudy (anonymous):

I figured it out!

OpenStudy (anonymous):

Good for you!!!

OpenStudy (anonymous):

perimeter: 10^2 + 24^2 =c^2 100+576=c^2 c=26 26+24+10 p=60 ft.

OpenStudy (anonymous):

oh ok then. :)

OpenStudy (anonymous):

The area is A = .5(b*h) A = .5(10ft * 24ft) A = 120ft^2 is this correct!